For the following linear system, find the values of x1 and x2 that satisfy this system using the LU factorization technique. Please show all your work in details. Failing to show your work in details might lead you to zero. 4x1 - 8x2 = 40 ,    -x1 + 6x2 = 50
Answer #1

480 - 8x₂ 40 2 -X+ 6 x 2 = 50 by 4 and multiply equation ④ in equation 0 add > 4x - 8x₂ 40 200 4X2 + 24X2 + 1642 = 240 X2 240Now by 3 multiply Squation o 14 and then by and equation ③ add both 120 12 xq 24%, z 200 - 4x + 24x₂ 320 as exy + 0 = 320